In youngs double slit experiment, the slits are 0.5 mm apart and interference is observed on a screen placed at a distance of 100 cm from the slits. It is found that 9 th bright fringe is at a distance 8.835 mm from the second dark fringe on the same side of the centre of the fringe pattern. Find the wavelength of light used?
1
Expert's answer
2019-09-02T09:18:59-0400

Let us use notation: "d = 0.5 mm" - distance between the slits, "D = 1m" - distance from the slits to the screen, "l_9" - distance from the center of the pattern to 9th maximum, "l_2" - distance from the center of the pattern to 2nd minimum.

The condition for the mth maximum is "d \sin \theta = m \lambda" , and for nth minimum is "d \sin \theta = (n + \frac{1}{2}) \lambda", where "\lambda" is the wavelength. Since the distance from the slit to the screen is quite big, let us use approximation "\sin \theta \approx \tan \theta = \frac{y}{D}" , where "y" is the distance from the center of the pattern to particular maximum/minimum. Hence, one has equations "\frac{d l_9}{D} = 9 \lambda" , "\frac{d l_2}{D} = (2+ \frac{1}{2}) \lambda" for maximum and minimum respectively. Subtracting the second equation from the first, obtain "\frac{d}{D}(l_9 - l_2) = 7.5 \lambda". The distance between the maximum and minimum is given, "l_9 - l_2 = 8.835 mm", therefore the wavelength is "\lambda = \frac{2 d (l_9 - l_2)} {15 D} = 589 nm".